# From affine algebraic racks to Leibniz algebras

### v1

@@ Abstract.
We introduce analogues of algebraic groups called algebraic racks, which are pointed rack objects in the category of schemes over a field $k$. Addressing a problem of Loday, we construct a functor assigning a canonical Leibniz algebra to every affine algebraic rack. This functor recovers the Lie algebras of affine algebraic groups (via conjugation quandles) and the Leibniz algebras of algebraic Lie racks, and it is compatible with closed subracks and left ideals. We also compare properties of affine algebraic racks, commutative corack algebras, and their Leibniz algebras.
@@
Affine algebraic group, algebraic rack, corack algebra, Leibniz algebra, Lie algebra, pointed rack object, quandle
@@ 1. Introduction
This paper introduces algebraic racks, which are pointed racks internal to the category of schemes of finite type over a field $k$. We show that affine algebraic racks provide a solution to a version of the well-known coquecigrue problem for Leibniz algebras over arbitrary fields $k$.
@@
Loday’s [loday] original version of the coquecigrue problem from 1993 seeks algebraic structures on smooth manifolds that differentiate to Leibniz algebras in a way that recovers the Lie group–Lie algebra correspondence. Leibniz algebras are vector spaces $\mathfrak{q}$ equipped with a bilinear mapping $\cdot,\cdot]\colon\mathfrak{q}\times\mathfrak{q}\to\mathfrak{q}$ called the Leibniz bracket that satisfies the derivation form of the Jacobi identity
$[X,[Y,Z]]=[[X,Y],Z]+[Y,[X,Z]].$
Bloh [bloh] introduced Leibniz algebras in 1965 to generalize Lie algebras, which are precisely those Leibniz algebras whose Leibniz bracket is alternating. See [leibniz] for an introduction to Leibniz algebras.
@@
1.1. Main theorem
Kinyon also posed an analogue of the coquecigrue problem over arbitrary fields $k$ that replaces Lie groups with affine algebraic groups. Save for unpublished work of Bašić and Škoda [skoda] in 2008 in the case that $\operatorname{char}(k)=0$, the author is unaware of any literature in this direction. To address this, we prove the following. Let ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})$ and $\mathsf{Leib}_{k}$} denote the categories of affine algebraic racks over $k$ and Leibniz algebras over $k$, respectively.
@@ 1.1. Main theorem
Let $Q=\operatorname{Spec}(A)$ be an affine algebraic rack with unit $e\in Q$, and let $\mathfrak{q}\coloneq\operatorname{Der}_{k}(A,k)$ be the vector space of $k$-derivations of $A$. Define a Leibniz bracket $\cdot,\cdot]\colon\mathfrak{q}\times\mathfrak{q}\to\mathfrak{q}$ by convolution with respect to the corack operation $\nabla$ of $A$ (see Section 4.1):
$D,E]\coloneq(D\otimes E)\circ\nabla.$
Then the assignment $Q\mapsto(\mathfrak{q},[\cdot,\cdot])$ defines a covariant functor ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})\to{\mathsf{Leib}_{k}}$ sending morphisms of affine algebraic racks $\varphi\colon Q\to R$ to their derivatives $(d\varphi)_{e}\colon\mathfrak{q}\to\mathfrak{r}$ at $e$.
Furthermore, if $G$ is an affine algebraic group and $Q=\operatorname{Conj}(G)$ is its conjugation quandle scheme, then $(\mathfrak{q},[\cdot,\cdot])$ is the Lie algebra of $G$. On the other hand, if $k=\mathbb{R}$ or $\mathbb{C}$ and $Q(k)$ is a Lie rack, then $(\mathfrak{q},[\cdot,\cdot])$ is the Leibniz algebra of $Q(k)$ as a Lie rack.
@@
We prove the theorem piecemeal in Propositions 6.2, 6.6, and 6.8 and Remark 6.3. ∎
@@
Theorem 1.1 is an algebraic analogue of a result of Kinyon’s result for Lie racks in [kinyon]*Thm. 3.4. Compared to that result, two advantages of Theorem 1.1 are that it holds for all fields $k$ and gives an explicit formula for the Leibniz bracket $\cdot$$\cdot$] that does not require computing any derivatives. That said, the Leibniz bracket agrees with the derivative of the adjoint representation of $Q$ at the identity; see Proposition 6.2.
@@ 1.2. Literature review
While there does not appear to be any published work on the algebraic coquecigrue problem over arbitrary fields kk—see [skoda] for a preliminary report of unpublished work of Bašić and Škoda from 2008 in the case that $\operatorname{char}(k)=0$—there is a rich literature on the coquecigrue problem for smooth manifolds over $\mathbb{R}$ and $\mathbb{C}$. See [ongay] for a discussion of the literature up to 2014.
Although the first work on the coquecigrue problem is due to Datuashvili [Datuashvili] in 2004, most modern approaches to the problem employ Lie racks, which Kinyon [kinyon] introduced in 2007 in response to the problem. For example, Covez [covez] in 2013 introduced local augmented Lie racks, which address the coquecigrue problem with the caveat that an algebraic structure is only defined in a neighborhood of the identity. By contrast, Bordemann and Wagemann [bordemann] in 2016 introduced a global rack-theoretic solution with the caveat of not being functorial.
@@
Notation and conventions.
We work over a field $k$ of arbitrary characteristic. By a scheme, we mean a scheme of finite type over $k$. Unless we are specifically discussing Leibniz algebras or Lie algebras, by a $k$-algebra we mean a finitely generated commutative associative $k$-algebra. Let $\mathsf{Sch}_{k}$}, $\mathsf{Aff}_{k}$}, $\mathsf{CAlg}_{k}$}, $\mathsf{Leib}_{k}$}, $\mathsf{Grp}$, $\mathsf{Grp}({\mathsf{Sch}_{k}})$, and $\mathsf{Set}$ denote the categories of schemes, affine schemes, $k$-algebras, Leibniz algebras over $k$, groups, algebraic groups (that is, group objects in $\mathsf{Sch}_{k}$}), and sets, respectively. Denote the $k$-algebra of dual numbers by $k[\delta]\coloneq k[\delta]/(\delta^{2})$. Given a morphism of affine schemes $\varphi\colon\operatorname{Spec}(B)\to\operatorname{Spec}(A)$, let $\varphi^{\sharp}$ denote the induced homomorphism $A\to B$.
@@ 1.3. Organization of the paper
In Section 2, we recall the definition of the category of pointed racks $\mathsf{Rack}_{\ast}$} and introduce pointed rack objects internal to cartesian monoidal categories.
In Section 3, we introduce the categories of algebraic racks ${\mathsf{Rack}_{\ast}}({\mathsf{Sch}_{k}})$ and corack algebras ${\mathsf{Corack}_{\ast}}({\mathsf{CAlg}_{k}})$. We also discuss the functor of points approach for algebraic racks.
In Section 4, we specialize our attention to the category of affine algebraic racks ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})$, which is equivalent to the opposite category $({\mathsf{Corack}_{\ast}}({\mathsf{CAlg}_{k}}))^{\mathrm{op}}$. As examples, we compute the corack algebras corresponding to three classes of affine algebraic racks. We also show that corack algebras are almost never coassociative or cocommutative, unlike Hopf algebras of algebraic groups.
In Section 5, we discuss our approach to proving Theorem 1.1, and we recall various facts about $k$-derivations and tangent spaces.
In Section 6, we prove Theorem 1.1. A major step in the proof is to show that the Leibniz bracket $\cdot$$\cdot$] in Theorem 1.1 agrees with the scheme-theoretic derivative $\operatorname{ad}\coloneq(d\operatorname{Ad})_{e}$ of the adjoint representation of $Q$ at the identity $e$.
In Section 7, we illustrate Theorem 1.1 with a class of affine algebraic racks $\operatorname{OL}_{n}$ whose Leibniz algebras $\mathfrak{ol}_{n}(k)$ are not generally Lie algebras. Then we provide several results relating the structure of an affine algebraic rack to that of its Leibniz algebra, including sufficient conditions for the latter to be abelian. These results recover the analogous results for Lie algebras of affine algebraic groups.
@@ Acknowledgments.
An early draft of this paper was submitted as the final project for a course on affine algebraic groups taught by Prof. Bogdan Ion at the University of Pittsburgh in fall 2025.
2. Pointed rack objects
@@ 2.1. Pointed racks
To motivate the definition of a pointed rack object, we begin in the set-theoretic setting of ordinary rack and quandle theory. Recall (from, say, [kinyon]) that a pointed rack consists of a set $Q$, a distinguished element $e\in Q$ called the unit or identity, and two binary operations $\triangleright,\triangleright^{-1}\colon Q\times Q\to Q$ satisfying the following five axioms [^1] Some authors swap the roles of $\triangleright$ and $\triangleright^{-1}$ in the definition of a rack. Of course, this choice is purely notational. This paper adopts the convention that $\triangleright$ is left-distributive because this seems to be the more common convention in the literature on Lie racks and quandle varieties. for all elements $x,y,z\in Q$:
@@
If $x\triangleright(x\triangleright y)=y$ (equivalently, $x\triangleright y=y\triangleright^{-1}x$) for all $x,y\in Q$, then we say that $Q$ is involutory. If $x\triangleright x=x$ for all $x\in Q$, then we say that $Q$ is idempotent or a (pointed) quandle. If $Q$ and $R$ are pointed racks, then a function $\varphi\colon Q\to R$ is called a pointed rack homomorphism if
@@
for all $x,y\in Q$. (In this case, $\varphi$ also preserves the right-distributive operations $\triangleright_{Q}^{-1}$ and $\triangleright_{R}^{-1}$; this is easiest to check using the notation in Section 2.1.2.) Let $\mathsf{Rack}_{\ast}$} be the category of pointed racks.
2.1.1.
@@
2.1.2.
Before we give an example of a pointed rack that is not a quandle, we note that in order to construct a pointed rack $Q$ with identity element $e$, it is enough to define a binary operation $\triangleright\colon Q\times Q\to Q$ satisfying axioms 2, 4, and 5 such that for all $x\in Q$, the function $L_{x}\colon Q\to Q$ defined by $y\mapsto x\triangleright y$ is bijective. This data completely determines the right-distributive binary operation $\triangleright^{-1}\colon Q\times Q\to Q$. In fact, $L_{x}$ is an automorphism of $Q$ called left multiplication by $x$; its inverse $L_{x}^{-1}$ is given by $y\mapsto y\triangleright^{-1}x$. See [fenn] for further discussion.
2.1.3.
@@
2.1.4.
Finally, we recall from [ryder] that if $Q$ is a pointed rack, then a subset $\{e\}\subseteq R\subseteq Q$ is called a (pointed) subrack if $L_{r}(R)=R$ for all $r\in R$ and a left ideal [^2] One may similarly define a right ideal of $Q$ by requiring that $r\triangleright x\in R$ for all $r\in R$ and $x\in Q$. However, this condition implies that $R=Q$; see [elhamdadi]*Prop. 4.6. or normal subrack if $L_{x}(R)=R$ for all $x\in Q$. [^3] Some papers, such as [elhamdadi], call our notion of a left ideal a “right ideal.” We deviate from this terminology to more closely analogize the definition of left ideals in rings. Moreover, [elhamdadi] and some other papers only require that $L_{r}(R)\subseteq R$ and $L_{x}(R)\subseteq R$ in their definitions of subracks and left ideals. These weaker conditions do not suffice to make $R$ a rack in general; see [kamada]*Sec. 2. In particular, left ideals are also pointed subracks; the converse is not true in general.
@@
2.2. Pointed rack objects

### v2

@@ Abstract.
We introduce analogues of algebraic groups called algebraic racks, which are pointed rack objects in the category of schemes over a ground field. Addressing a problem of Loday [loday], we construct functors assigning left and right Leibniz algebras to affine algebraic racks. These functors are compatible with closed subracks and ideals, and they recover the Lie algebras of linear algebraic groups (via conjugation quandles) and the Leibniz algebras of algebraic Lie racks. We also study properties of coordinate algebras and Leibniz algebras of affine algebraic racks. Finally, we use rack schemes to functorially construct (co-)nondegenerate Yang–Baxter operators in various categories.
@@
Affine algebraic rack, algebraic group, braid equation, conjugation quandle, Leibniz algebra, nondegenerate Yang–Baxter solution, pointed rack object
@@ 1. Introduction
This paper introduces analogues of algebraic groups called rack schemes (resp. algebraic racks), which are racks (resp. pointed racks) internal to the category of schemes of finite type over a ground field $k$. We show that affine algebraic racks provide a solution to a version of the well-known coquecigrue problem for left and right Leibniz algebras over arbitrary fields $k$ (see Theorem 1.1). We also use rack schemes to construct solutions to the Yang–Baxter equation in the categories of schemes, commutative $k$-algebras, and sets (see Sections 12 and 13.2).
@@
Loday’s [loday] original version of the coquecigrue problem from 1993 seeks algebraic structures on smooth manifolds that differentiate to Leibniz algebras in a way that recovers the Lie group–Lie algebra correspondence. Bloh [bloh] introduced Leibniz algebras in 1965 to generalize Lie algebras.
@@
1.1. Main results
Kinyon [kinyon] posed an analogue of the coquecigrue problem over arbitrary fields $k$ that replaces Lie groups with algebraic groups. Save for a preliminary report of Bašić and Škoda [skoda] from 2008, the author is unaware of any work in this direction. To address this, we prove the following theorem. Working over a ground field $k$, let ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})$, $\mathsf{Leib}_{k}$}, and $\mathsf{RLeib}_{k}$} denote the categories of affine algebraic racks, left Leibniz algebras, and right Leibniz algebras, respectively.
@@ 1.1. Main results
Let $Q=\operatorname{Spec}(A)$ be an affine algebraic rack with unit $e\in Q$, and let $\mathfrak{q}\coloneq\operatorname{Der}_{k}(A,k)$ be the vector space of $k$-derivations of $A$. Define two Leibniz brackets $\cdot,\cdot],\{\cdot,\cdot\}\colon\mathfrak{q}^{2}\to\mathfrak{q}$ by convolution with respect to the corack operations $(\nabla,\nabla^{-1})\coloneq(\triangleright^{\sharp},(\triangleright^{-1})^{\sharp})$ of $A$ (see Section 6.1):
$D,E]\coloneq(D\otimes E)\circ\nabla,\qquad\{D,E\}\coloneq(D\otimes E)\circ\nabla^{-1}.$
Then the assignments $Q\mapsto(\mathfrak{q},[\cdot,\cdot])$ (resp. $Q\mapsto(\mathfrak{q},\{\cdot,\cdot\})$) and $\varphi\mapsto(d\varphi)_{e}$ define a covariant functor ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})\to{\mathsf{Leib}_{k}}$ (resp. ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})\to{\mathsf{RLeib}_{k}}$).
Furthermore, if $Q=\operatorname{Conj}(G)$ is the conjugation quandle scheme of an affine algebraic group $G$, then $(\mathfrak{q},[\cdot,\cdot])$ is the Lie algebra of $G$. On the other hand, if $k=\mathbb{R}$ or $\mathbb{C}$ and $Q(k)$ is a Lie rack, then $(\mathfrak{q},[\cdot,\cdot])$ is the Leibniz algebra of $Q(k)$ as a Lie rack.
@@
For $\cdot$$\cdot$], we prove the theorem piecemeal in Propositions 8.2, 8.6, and 8.7 and Remark 8.3. The proof for $\{\cdot,\cdot\}$ is similar. ∎
@@
Theorem 1.1 is an algebraic analogue of a result of Kinyon’s result for Lie racks in [kinyon]*Thm. 3.4. Compared to that result, two advantages of Theorem 1.1 are that it holds for all fields $k$, endows $\mathfrak{q}$ with both left and right Leibniz brackets, and gives explicit formulas for both brackets. Like in Kinyon’s construction, the left Leibniz bracket $\cdot$$\cdot$] agrees with the derivative of the adjoint representation of $Q$ at the identity in the sense of Proposition 8.2.
@@
1.1.2.
As an application of rack schemes over commutative rings $S$, we also construct solutions to the Yang–Baxter equations in various categories; see Sections 12 and 13.2. Our constructions employ a fully faithful functor from the category of rack schemes to the category of Yang–Baxter schemes (that is, braided objects in the category of schemes over $S$); see Theorem 11.1.
[…]
@@ 1.2. Literature review
It appears that the only existing work on the algebraic coquecigrue problem over arbitrary fields $k$ is a preliminary report of unpublished work of Bašić and Škoda [skoda] from 2008 in the case that $\operatorname{char}(k)=0$. On the other hand, there is a rich literature on the coquecigrue problem for smooth manifolds over $\mathbb{R}$ and $\mathbb{C}$. See [ongay] for a discussion of the literature up to 2014.
Although the first work on the coquecigrue problem is due to Datuashvili [Datuashvili] in 2004, most modern approaches to the problem employ Lie racks, which Kinyon [kinyon] introduced in 2007 in response to the problem. For example, Covez [covez] in 2013 used local augmented Lie racks to provide an analogue of Lie’s third theorem for Leibniz algebras over $\mathbb{R}$ with the caveat that a rack structure is only defined in a neighborhood of the identity. By contrast, Bordemann and Wagemann [bordemann] in 2017 introduced a non-functorial way to integrate Leibniz algebras over $\mathbb{R}$ into Lie racks.
@@
1.2.2.
[…]
1.3. Notation and conventions
Let $k$ be a field, and let $S$ be a commutative ring. By a scheme, we mean a scheme of finite type over $k$ or $S$. Unless we are specifically discussing Leibniz algebras, by $k$-algebras we mean finitely generated commutative associative algebras over $k$, and similarly for $S$-algebras. Denote the $k$-algebra of dual numbers by $k[\delta]\coloneq k[\delta]/(\delta^{2})$. Given a morphism of affine schemes $\varphi\colon\operatorname{Spec}(B)\to\operatorname{Spec}(A)$, let $\varphi^{\sharp}\colon A\to B$ denote the induced ring homomorphism.
Let $\mathsf{Sch}_{S}$}, $\mathsf{Aff}_{S}$}, $\mathsf{CAlg}_{S}$}, $\mathsf{Leib}_{k}$}, $\mathsf{RLeib}_{k}$}, $\mathsf{Grp}$, $\mathsf{Grp}({\mathsf{Sch}_{k}})$, $\mathsf{Set}$, $\mathsf{Rack}$, and $\mathsf{Rack}_{\ast}$} denote the categories of schemes over $S$, affine schemes over $S$, $S$-algebras, left Leibniz algebras over $k$, right Leibniz algebras over $k$, groups, algebraic groups over $k$, sets, racks, and pointed racks, respectively.
Given an object $X$ in a monoidal category $(\mathcal{C},\otimes)$, let $X^{n}\coloneq\bigotimes^{n}_{i=1}X$ denote the $n$-fold tensor product of $X$. If $(\mathcal{C},\otimes)$ is braided monoidal, let $\tau$ denote the braiding $\tau\colon X^{2}\xrightarrow{\sim}X^{2}$. If $(\mathcal{C},\times)$ is cartesian monoidal, then let $\pi_{1},\pi_{2}\colon X^{2}\to X$ denote the projections (so, for example, $\tau=(\pi_{2},\pi_{1})$).
@@ 1.4. Organization of the paper
The first few sections of this paper are dedicated to preliminaries. In Section 2, we recall the definitions of Leibniz algebras and pointed racks and give several important examples.
In Section 3, we introduce pointed rack objects internal to cartesian monoidal categories.
In Section 4, we introduce the categories of algebraic racks ${\mathsf{Rack}_{\ast}}({\mathsf{Sch}_{k}})$ and affine algebraic racks ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})$. We also discuss the functor of points approach for algebraic racks.
In Section 5, we introduce the category of (commutative) corack algebras ${\mathsf{Corack}_{\ast}}({\mathsf{CAlg}_{k}})$.
1.4.1.
Looking toward Theorem 1.1, we specialize our attention to the category of affine algebraic racks ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})$ over a ground field $k$. In Section 6, we discuss the anti-equivalence between ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})$ and ${\mathsf{Corack}_{\ast}}({\mathsf{CAlg}_{k}})$. As examples, we compute the corack algebras corresponding to three classes of affine algebraic racks. We also show that, unlike Hopf algebras of algebraic groups, corack algebras are coassociative or cocommutative only in trivial cases.
In Section 7, we discuss our approach to proving Theorem 1.1. In particular, we define adjoint representations $\operatorname{Ad}$, $\operatorname{ad}$ and dual adjoint representations $\operatorname{Ad}^{-1},$ $\operatorname{ad}^{-1}$ of affine algebraic racks and their tangent spaces. We also recall various facts about $k$-derivations and tangent spaces.
In Section 8, we prove Theorem 1.1. A major step in the proof is to show that, in a certain sense, the Leibniz brackets $\cdot$$\cdot$] and $\{\cdot,\cdot\}$ respectively agree with the maps $\operatorname{ad}$ and $\operatorname{ad}^{-1}$.
In Section 9, we provide several results relating the structure of an affine algebraic rack to that of its Leibniz algebra, including sufficient conditions for the latter to be abelian. These results recover the analogous results for Lie algebras of affine algebraic groups.
1.4.2.
Next, we discuss how rack schemes $\mathsf{Rack}({\mathsf{Sch}_{S}})$ over a commutative ring $S$ induce solutions to the Yang–Baxter equation in $({\mathsf{Sch}_{S}},\times_{S})$, $(\mathsf{Set},\times)$, and $({\mathsf{CAlg}_{S}},\otimes_{S})$. In Section 10, we recall definitions relating to Yang–Baxter operators $\mathsf{YB}(\mathcal{C})$ internal to a tensor category $(\mathcal{C},\otimes)$.
[…]
In Section 12, we apply the results of Sections 10 and 11 to $\mathsf{Sch}_{S}$} and $\mathsf{Aff}_{S}$}. In particular, we discuss how rack schemes induce nondegenerate Yang–Baxter operators in $\mathsf{Sch}_{S}$} and $\mathsf{Set}$ and co-nondegenerate Yang–Baxter operators in $\mathsf{CAlg}_{S}$}.
1.4.3.
In Section 13, we illustrate Theorem 1.1 with a class of affine algebraic racks $\operatorname{OL}_{n}$ whose Leibniz algebras $\mathfrak{ol}_{n}(k)$ are not generally Lie algebras. To illustrate the constructions in Section 12, we also compute the induced Yang–Baxter operators in $\mathsf{Aff}_{k}$} and $\mathsf{CAlg}_{k}$}.
@@ Acknowledgments.
An early draft of this paper was submitted as the final project for a course on affine algebraic groups taught by Prof. Bogdan Ion at the University of Pittsburgh in the fall 2025 semester. I thank Prof. Ion for helpful comments and suggestions.
2. Preliminaries
@@
2.1. Leibniz algebras
We recall the definitions of left and right Leibniz algebras over $k$.
Let $\mathfrak{q}$ be a vector space over $k$, and let $\cdot,\cdot],\{\cdot,\cdot\}\colon\mathfrak{q}^{2}\to\mathfrak{q}$ be bilinear mappings. We call $\cdot$$\cdot$] a (left) Leibniz bracket and $(\mathfrak{q},[\cdot,\cdot])$ a left Leibniz algebra if, for all $X\in\mathfrak{q}$, the map $\operatorname{ad}_{X}\colon\mathfrak{q}\to\mathfrak{q}$ defined by $Y\mapsto[X,Y]$ is a derivation of $\cdot$$\cdot$]. That is, the left derivation form of the Jacobi identity
$[X,[Y,Z]]=[[X,Y],Z]+[Y,[X,Z]]$
holds for all $X,Y,Z\in\mathfrak{q}$.
Similarly, we call $\{\cdot,\cdot\}$ a (right) Leibniz bracket and $(\mathfrak{q},\{\cdot,\cdot\})$ a right Leibniz algebra if for all $X\in\mathfrak{q}$, the map $\operatorname{ad}_{X}^{-1}\colon\mathfrak{q}\to\mathfrak{q}$ defined by $Y\mapsto\{Y,X\}$ is a derivation of $\{\cdot,\cdot\}$:
$\{\{X,Y\},Z\}=\{\{X,Z\},Y\}+\{X,\{Y,Z\}\}.$
If $(\mathfrak{q},[\cdot,\cdot]_{\mathfrak{q}})$ and $(\mathfrak{r},[\cdot,\cdot]_{\mathfrak{r}})$ are left Leibniz algebras, then a $k$-linear map $\varphi\colon\mathfrak{q}\to\mathfrak{r}$ is called a (left) Leibniz algebra homomorphism if it preserves the respective Leibniz brackets. That is,
$\varphi([X,Y]_{\mathfrak{q}})=[\varphi(X),\varphi(Y)]_{\mathfrak{r}}$
for all $X,Y\in\mathfrak{q}$. Right Leibniz algebra homomorphisms are defined identically.
2.1.1.
For example, a Lie algebra is precisely a left or right Leibniz algebra whose Leibniz bracket is alternating; that is, $[X,X]=0$ for all $X\in\mathfrak{q}$.
There are many examples of left Leibniz algebras that are not Lie algebras. Of particular interest in this paper is the following construction of Weinstein [weinstein] in 2000; see [omni]. Given a vector space $V$, consider the vector space $\mathfrak{gl}(V)\oplus V$. (This is the underlying vector space of the Lie algebra $\mathfrak{aff}(V)$ of the linear algebraic group $\mathrm{Aff}(V)$ of affine transformations of $V$.)
Define a left Leibniz bracket on $\mathfrak{gl}(V)\oplus V$ by
(2.1) $(X,v),(Y,w)]\coloneq(XY-YX,Xw).$
Then $\mathfrak{ol}(V)\coloneq(\mathfrak{gl}(V)\oplus V,[\cdot,\cdot])$ is a hemisemidirect product left Leibniz algebra (see [kinyon2]) called an omni-Lie algebra. Despite its name, $\mathfrak{ol}(V)$ is a Lie algebra if and only if $\dim(V)=0$. (See [weinstein] for the original motivation for this name; cf. [omni]*Prop. 3.1.) If $V=k^{n}$, then we denote $\mathfrak{ol}_{n}(k)\coloneq\mathfrak{ol}(k^{n})$.
@@ 2.2. Pointed racks
To motivate the definition of a pointed rack object, we begin in the set-theoretic setting of ordinary rack and quandle theory. Recall (from, say, [kinyon]) that a pointed rack consists of a set $Q$, a distinguished element $e\in Q$ called the unit or identity, and two binary operations $\triangleright,\triangleright^{-1}\colon Q^{2}\to Q$ satisfying the following five axioms [^1] Some authors swap the roles of $\triangleright$ and $\triangleright^{-1}$ in the definition of a rack. Of course, this choice is purely notational. This paper adopts the convention that $\triangleright$ is left-distributive because this seems to be the more common convention in the literature on Lie racks and quandle varieties. for all elements $x,y,z\in Q$:
@@
If $x\triangleright(x\triangleright y)=y$ (equivalently, $x\triangleright y=y\triangleright^{-1}x$) for all $x,y\in Q$, then we say that $Q$ is involutory. If $x\triangleright x=x$ for all $x\in Q$, then we say that $Q$ is idempotent or a (pointed) quandle.
Given pointed racks $Q$ and $R$, a function $\varphi\colon Q\to R$ is called a pointed rack homomorphism if
@@
for all $x,y\in Q$. (In this case, $\varphi$ also preserves the right-distributive operations $\triangleright_{Q}^{-1}$ and $\triangleright_{R}^{-1}$. Verifying this fact is easiest when using the notation $L_{x}^{\pm 1}$ in Section 2.2.2.) We can similarly define racks and rack homomorphisms by omitting every axiom that involves the unit $e$; cf. [fenn].
Note that every pointed rack $Q$ also satisfies the following identities for all $x\in Q$:
  1. (Q4’)
    (Inverse fixing) $x\triangleright^{-1}e=x$.
  2. (Q5’)
    (Inverse fixedness) $e\triangleright^{-1}x=e$.
2.2.1.
@@
2.2.2.
Before we give an example of a pointed rack that is not a quandle, we note that in order to construct a pointed rack $Q$ with identity element $e$, it is enough to define a binary operation $\triangleright\colon Q^{2}\to Q$ satisfying axioms 2, 4, and 5 such that for all $x\in Q$, the function $L_{x}\colon Q\to Q$ defined by $y\mapsto x\triangleright y$ is bijective. This data completely determines the right-distributive binary operation $\triangleright^{-1}\colon Q^{2}\to Q$.
In fact, for all elements $x\in Q$, the function $L_{x}$ is a rack automorphism of $Q$ called left multiplication by $x$. The inverse function $L_{x}^{-1}$ is given by $y\mapsto y\triangleright^{-1}x$. See [fenn] for further discussion.
2.2.3.
@@
2.2.4.
Finally, we recall from [ryder] that if $Q$ is a pointed rack, then a subset $\{e\}\subseteq R\subseteq Q$ is called a (pointed) subrack if $L_{r}(R)=R$ for all $r\in R$ and a left ideal [^2] One may similarly define a right ideal of $Q$ by requiring that $r\triangleright x\in R$ for all $r\in R$ and $x\in Q$. However, as observed in [elhamdadi]*Prop. 4.6, this condition implies that $R=Q$. Indeed, $e\in R$, so $x=e\triangleright x\in R$ for all $x\in Q$. or normal subrack if $L_{x}(R)=R$ for all $x\in Q$. [^3] Some papers, such as [elhamdadi], call our notion of a left ideal a “right ideal.” We deviate from this terminology to more closely analogize the definition of left ideals in rings. Moreover, [elhamdadi] and some other papers only require that $L_{r}(R)\subseteq R$ and $L_{x}(R)\subseteq R$ in their definitions of subracks and left ideals. These weaker conditions do not always guarantee that $R$ a rack; see [kamada]*Sec. 2 for counterexamples. In particular, left ideals are also pointed subracks; the converse is not true in general.
@@
3. Pointed rack objects
[…]
3.1.